# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v2

@@ 4.8. Interpretability Analysis
4.9. Limitation and Future Prospects
TEFN is a naive and unmodified model. We lack skills in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method. Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well. In addition, for the fusion of mass distribution in the face of the failure of the classic Dempster Shafer rule for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern. In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels. In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution. Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic. Effective parameter initialization may exist, for example, initializing corresponding parameters in a non neural network way may lead to faster convergence.
5. Conclusion
This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets. TEFN treats different channels and time points as distinct information sources, using evidence theory to construct basic probability assignments and discard irrelevant data. This selective information fusion significantly reduces the computational burden, enabling TEFN to efficiently process and predict within acceptable time frames. The model’s parameter efficiency, robustness, and interpretability make it a valuable tool for time series forecasting tasks involving large and complex datasets.
Acknowledgements.
The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078)

### v3

@@ 4.9 Interpretability Analysis
4.10 Limitation and Future Prospects
TEFN is a naive and unmodified model. There is not engineering experience in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method. Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well. In addition, for the fusion of mass distribution in the face of the failure of the classic DSR for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern. In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels. In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution. Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic. Effective parameter initialization may exist, for example, initializing corresponding parameters in a non-neural network way may lead to faster convergence.
5 Conclusion
This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets. TEFN treats different channels and time points as distinct information sources, using evidence theory to construct basic probability assignments and discard irrelevant data. This selective information fusion significantly reduces the computational burden, enabling TEFN to process and predict within acceptable time frames efficiently. The model’s parameter efficiency, robustness, and interpretability make it a valuable tool for time series forecasting tasks involving large and complex datasets.